Sovle the pairs of equations using the substitituion method.
x - y = 26
3x - 8y = 3
So far I have: 3(26 + y) - 8y = 3 69y - 8y = 3

Answer by fractalier(6550)

You're okay to here
3(26 + y) - 8y = 3
now distribute and get
78 + 3y - 8y = 3
78 - 5y = 3
-5y = -75
y = 15 and then
x = 41